On average, each person in the industrialized world is responsible for the emission of 10,000 kg of carbon dioxide () every year. This includes that you generate directly, by burning fossil fuels to operate your car or your furnace, as well as generated on your behalf by electric generating stations and manufacturing plants. is a greenhouse gas that contributes to global warming. If you were to store your yearly emissions in a cube at STP, how long would each edge of the cube be?
Answer #1
Carbon dioxide has a molar mass of 44.01g/mol
So, 10,000kg would be:

10,000,000g * (1mol/44.01g) = 227,221mol

by the ideal gas law, at 1atm and 273K (STP), this would have volume:
PV = nRT
(1atm)V = (227221mol)(0.0821L*atm/mol*K)(273K)

V = 5,092,774L or 5,092.774m^3

If this was contained in a cube,
V = s^3
5092.774m^3 = s^3
s = 17.2m

each side would measure 17.2m